OpenStudy (anonymous):

Find the x-intercepts of the parabola with vertex (6,-5) and y-intercepts (0,175). Write the answer in this form: (x1,y1),(x2,y2).

OpenStudy (anonymous):

help? :(

OpenStudy (anonymous):

1) (y - k)^2 = 4a(x - h) 2) (y- -5)^2 = 4a(x-6) 3) (175+5)^2=4a(0-6) 4) (180)^2 = 4a (-6) 5) 32400 = 4a(-6) 6) -5400 = 4a 7) -1350=a 8)(y +5)^2 = 4 * -1350(x - 6) at this point it is easier to graph it on a calculator set it as : y= [+/- square root of -5400(x - 6)] -5 (b/c it could be either pos. or neg.)

OpenStudy (anonymous):

w/o a calc : 5^2 = -5400(x-6) .....remember y=0 at an x intercept 25/-5400= x-6 -.0046=x-6 5.9954=x